Grande03:

Hamid published a book of poems. He sold 35 copies in the first week for $10 per copy. He wanted to increase the price of each book to increase his revenue. He found out that for every 50-cent increase, x, in the price per copy, the average number of books he sold decreased by 2 per week. Use this information to create expressions that model the number of books he sold and the revenue he got. The number of books sold is given by the expression ______, and the revenue is given by the expression ______.

Grande03:

helpp

mdobbs6856:

y = (10 + 0.5x)(35 − 2x) y = 350 − 20x + 17.5x − x^2 y = 350 − 2.5x − x^2 y = -x^2 − 2.5x + 350 MakE sense?

Grande03:

okay, but which one goes in which blank?

Grande03:

@mdobbs6856

mdobbs6856:

what do you think?

Grande03:

you literally gave me four equations, idk

mdobbs6856:

The first equation came from the verbal equation you put. Which was the original problem and the final answer was the last equation sooooo.....?

Grande03:

okay but there are two blanks not one

Grande03:

so which blank did yu answer?

mdobbs6856:

The number of books sold is given by the expression y = (10 + 0.5x)(35 − 2x), and the revenue is given by the expression y = -x^2 − 2.5x + 350.

Grande03:

that has to be wrong because it doesn't even give me the tool to put x^2 and it only doesn't give me tools when I don't need them

Grande03:

its incorrect

Grande03:

not even close... idk where you got that from

mdobbs6856:

You asked for help

Grande03: